Conversions: 39. 88 kg-b 34. 1 mg mcg ml 12 oz= 35, mL 36. 2ts 37. 2000 mg= mL 43. 1 qt- ml 38.6.2 L=
Answer #1

ANSWER.

Conversions.

34) 1 mg = 1000 mcg.

35) 12 oz.

          1 oz = 30 ml.

    So

       12 oz = 30 * 12

                 = 360 ml.

36) 2t.

       1t = 15 ml.

   So

      2t = 30 ml.

37) 2000 mg.

        1g = 1000 mg.

      So

       2000 mg = 2g.

38) 6.2L.

        1 L = 1000 ml.

      So

        6.2L = 6.2 * 1000

                = 6200 ml.

39) 88kg.

        1kg = 2.205 lb.

     So

       88kg = 194.04 lb.

40) 3/4 gr.

        1gm = 1000 mg.

       So

          3/4 gm = 750 mg.

41) 6g.

        1g = 1000000.

      So

         6g = 6000000.

42) 1/2 T

         1T = 15 ml.

      So

         1/2 T = 7.5 ml.

43) 1 liquid quart(qt).

        1qt = 946.4 ml